# Nearly forgotten results in development of physical cosmology

### v2

@@ Abstract
NICA is a tool to investigate the early Universe in laboratory. It would be reasonable to recall some critical issues in physical cosmology development. GR was created by A. Einstein in 1915. In 1917 Einstein proposed the first (static) cosmological model. Soon after the A. Eddington proved that the model is unstable therefore it can not be realizable in nature. In 1922 and 1924 A. A. Friedmann found non-stationary solutions for cosmological equations written in the framework of GR. In 1927 G. Lemaître obtained very similar results and, in addition, he derived the Hubble law (E. Hubble obtained this law from observations). Unfortunately, G. Lemaître published his paper in not very popular Belgium journal. In 1931 Lemaître proposed the first version of hot Universe model (he called it hypothesis of the primeval atom). In his book ”The Primeval Atom – An Essay on Cosmogony” Lemaitre predicted even a background radiation as a signature of his model. At the end of 1940s G. Gamow and his students proposed his hot universe model where he explained primordial nucleosythesis of helium at the initial stage of the Universe evolution. One of the properties of Lemaître – Gamow model was a prediction of CMB radiation with a temperature around a few K. It was recalled that the discovery of CMB radiation was done by T. Shmaonov in 1956 and his paper was published in 1957 (several years before Penzias and Wilson). In 1965, 1970 E. B. Gliner proposed vacuum like equation of matter which could correspond to exponential explosion of the Universe which was later called inflation. For decades in USSR, A. A. Friedmann’s cosmologocal non-stationary models were treated as purely mathematical results without cosmological and astronomical applications. On September 16, 1925 passed away untimely and it would be reasonable to remind today his great contribution in physical cosmology since the authors of the book on Friedmann’s life and his scientific activity wrote “Similarly to Copernicus who forced the Earth to move, Friedmann forced the Universe to expand”.
@@ 1 Introduction
General Relativity (GR) was created after intensive conversations between A. Einstein and D. Hilbert in November 1915 [1, 2, 3]. Using his new theory of gravity in November 1915 A. Einstein explained the problem of the anomaly of Mercury’s motion, discovered in the middle of the 19th century by Le Verrier and remained unexplained for several decades, and also considered the problem of deflecting a ray of light in a gravitational field, and it turned out that the angle of deviation in general relativity exceeds the angle of deviation in the Newtonian theory of gravity by two times [4] (see, English translation in [5]).
In 1917 A. Einstein found a static cosmological solution assuming that the Universe is homogeneous and isotropic [6] (see English translation in [7]). However, the assumptions adopted in the first cosmological paper were very useful and despite the fact that assumptions about the homogeinity and isotropy of the universe do not look very plausible, such assumptions turn out to be applicable to the theoretical description of astronomical observations and are still used in a large number of works on cosmology and are often called the cosmological principle. To create a static cosmological solution and to compensate gravitational attraction A. Einstein introduced so-called $\Lambda$- term which corresponds to repulsion. In many textbooks and articles the authors repeated after G. Gamow that ”he [A. Einstein] remarked that the introduction of the cosmological term was the biggest blunder he ever made in his life” [8] (see also a similar sentence in a popular article [9])). However, taking into account Gamow’s personality, in particular, his predilection for jokes and practical jokes, and the different social status of Gamow and Einstein, it seems very strange that Einstein chose Gamow as an interlocutor to whom Einstein would admit to the biggest blunder of his entire life (see also discussions in [10, 11]). More over, in 1998 it was discovered an accelerated expansion of the Universe and in order to explain this phenomenon, it is necessary to introduce a non-zero value of the lambda term (or some generalization of it, which since 1999 has been called dark energy) into the Einstein equations.
When discussing Einstein’s static cosmological model, it should be noted that it is hardly applicable to describe the behavior of the universe because as a famous Russian mathematician V. I. Arnold repeated only stable mathematical solutions exist in physical world. Analyzing properties of the static cosmological model in 1930 A. Eddington ”working in conjunction with Mr. G. C. McVittie” proved that the Einstein’s static cosmological solution is unstable [12].
@@ 2 Start of GR studies in Russia
In the early 20th century, despite significant achievements in some areas of physical science development in Russia, world leaders in physics worked in Europe. Young Russians who wanted to acquire scientific skills at the cutting edge came to Europe for training and internship. One of these young people interested in science was a representative of the elite of the Russian empire (his father and grandfather were governors in Russian empire), Vsevolod Konstantinovich Frederiks (13.04.1885 (Warsaw) –- 06.01.1944 (Gorky)). He played a key role in initiating research in Russia in the field of GR [13, 14]. In 1903 Vsevolod Frederiks finished gymnasium in Nizhnij Novgorod and decided to go to Switzerland due to its convenient location to visit other European countries. In 1907 he finished Geneva University with specialisation in physics and in 1909 he got PhD degree in physics under supervision professor G. E. Guye (Geneve). Since 1911 Frederiks was an assistant under professor Woldemar Voigt [^2] In 1887 W. Voigt found transformations which were similar to Lorentz ones [15]. In addition, Voigt described group properties of his transformations as it was noted by H. Minkowski in his famous lecture delivered in 1908 [16] (see also discussions about a relation between Voigt and Lorentz transformations in [17, 18, 19, 20]). at the Theoretical division of Physics Institute at Göttingen. In 1914 the WWI began and Frederiks started to be a civil prisoner since he was a citizen of an enemy country. In 1910s famous mathematician David Hilbert was very interested in physics he was searching for educated physicist to discuss actual physical issues. Frederiks began working as Hilbert’s personal assistant, and Hilbert paid Frederiks a monetary reward from his personal funds. The management of the University of Göttingen did not approve this practice to contact with representative of an ebemt country in University and refused to allow Fredericks to enter the University in response to Hilbert’s appeal. In 1918 Frederiks came back in Russia and for a short period he worked in Institute of Physics and Biophysics in Moscow. In 1919 he moved to Petrograd and was a senior physicist at State Optical Institute, a member of Atomic commission, associate professor in Petrograd State University, professor in Pedagogical Institute, in 1920 he was a lecturer at Polytechnic Institute. In 1921 Frederiks published the first review on general relativity (GR) in Soviet Physics Uspekhi which was a leading physical journal in Russia [21]. In 1923 Frederiks was a senior physicist in Institute of Physics and Technology, the Institute was founded by A. F. Ioffe in 1918 and in 1924 it was named the Leningrad Institute of Physics and Technology (LIPT) or shortly the Ioffe Institute. V. K. Frederiks and A. A. Friedmann [^3] British expert in astronomy and history of science S. Mitton noted [22]: ”There are three ways to spell his surname: Fridman, Friedmann, and Friedman… Friedmann is the most common version in use today, largely in connection with cosmology. I [S. Mitton] use Friedman when referring to his publications because that is the version he used as author. A Bolshevik Government Decree issued 1917 December 23 eliminated double letters in Russian orthography. Typographers were required to enforce the reforms. The contemporary Russian literature invariably uses Friedman.” I [AZ] basically use Friedmann in the paper since this spelling was used in a majority of articles and books that I read before writing the paper including English translation of the remarkable book [23] written by the best experts in history of Russian cosmology. decided to wrote joint book on foundations of GR but in 1924 it was published only the first chapter of the book where tensor calculus was discussed (this project to write a joint book was not realized due to the untimely death of A. A. Friedmann). In Autumn 1927 Frederiks married Maria Dmitrievna Shostakovich (a sister of composer D. D. Shostakovich). In 1931 Frederiks was appointed as the Head of the Crystallization Laboratory in LPTI and in 1933 started to be the Head of anisotropic liquid Laboratory in LPTI. Frederiks studies in liquid crystal field were rather successful, in particular in 1927 V. Frederiks and A. Repiewa discovered the phase transition in liquid crystals under action of magnetic field. Now this property of liquid crystals is used in operation LC screens in different devices and it is called now the Frederiks transition [24]. In 1934 Frederiks got DSc degree without a formal defense, he was nominated by the LPTI Scientific Council as a candidate for corresponding member of Soviet Academy of Science but he was not elected. Due to very rapid development of contemporary physics in 1920s and 1930s it was necessary to write new books and V. K. Frederiks was the co-editor with A. . Afanasiev of Course on General Physics (I. K. Kikoin, Yu. B. Khariton were among authors of chapters in the book). Later, the textbook was republished where only A. . Afanasiev was indicated since Frederiks was in detention at this time. On October 20, 1936 V. K. Frederiks (among many other Leningrad scientists) was arrested as a defendant in the Pulkovo case. On May 23, 1937 Frederiks sentenced to 10 years in prison camps (Taishetlag). On January 6, 1944 Frederiks died. Only in 1957 his relatives got an official document on his death, where is the dash was written in the place of death. Lydia Chukovskaya (whose husband Matvey Bronstein [^4] M. Bronstein was one of the leaders of Soviet theoretical physics [25] and he was among the pioneers in quantum gravity research [26] (in the quoted paper C. Rovelli noted that an introduction of the graviton concept proposed by D. I. Blokhintsev and F. M. Galperin in 1934 was an important initial step in development of quantum gravity). was an another Pulkovo case defendant) received a similar document concerning a time and a place of the Bronstein death.
@@
There’s no doubt that Friedmann’s cosmological work laid the foundations of physical cosmology and is a source of pride for Russian and international science [27, 28]. For Friedmann’s centenary, which was celebrated in June 1988, a scientific biography was written [29] (this remarkable book was translated into English in 1993). In the book the authors wrote, ”He [Friedmann] discards a centuries-old tradition, which a priori, prior to all experience, considered the Universe eternal and forever motionless. He is making a real scientific revolution. Just as Copernicus forced the Earth to revolve around the Sun, so Friedmann forced the Universe to expand”. In the days of the Baldin seminar we celebrated a memory of this remarkable scientist. Alexander Alexandrovich Friedmann (Friedman) was born on 4(16).06.1888 in Sankt-Peterburg (SPb) and died on 16.09.1925 in Leningrad (Sankt-Peterburg (1703 – 1914), Petrograd (1914 – 1924), Leningrad (1924 – 1991), Sankt-Peterburg (1991 – till now)). In 1897 - 1906 Alexander Friedmann was a student at the Second SPb gymnasium, in 1905 he wrote the first mathematical paper (published in 1907). In 1906 - 1910 he was student at mathematical division of the faculty of physics and mathematics. In 1910 - 1913 he left in SPb University for a preparation for a professor position. In 1914 – 1916 Friedmann joined the Russian army as a volunteer, he served in aviation units. In 1918 – 1920 he was a professor of mechanics department of Perm University. In 1920 – 1924 he was a researcher at Atomic Commission in State Optical Institute. In 1920 – 1924 he was a professor at the Faculty of Physics and Mechanics of Petrograd Polytechnic Institute In 1920 – 1925 Friedmann was a senior physicist, head of mathematical bureau, scientific secretary and since February 1925 director of Main Geophysical Observatory In 1922 Friedmann published his first cosmological paper [30] (English translation of the paper was published in [31]). Soon after that Einstein published a short comment [32] (consisting a few sentences and one expression) where he claimed:”The results of the work on the non-stationary world contained in the mentioned work seem suspicious to me. In fact, it turns out that the solution indicated in it does not satisfy the field equations.” [^5] A Russian translation of the comment published in the special issue of Soviet Physics – Uspekhi dedicated to the 75th anniversary of Friedmann https://ufn.ru/ufn63/ufn63_7/Russian/r637g.pdf. Friedmann was extremely disappointed by this opinion of the GR creator. Friedmann consulted with Petrograd mathematicians and sent a lengthy letter to Einstein where Friedmann presented formulas confirming his conclusions. At the end of his letter Friedmann requested Einstein [23] ”…Should you find the calculations presented in my letter correct, please be so kind as to inform the editors of the Zeitschrift fur Physik about it; perhaps in this case you will publish a correction to your statement or provide an opportunity for a portion of this letter to be printed.” In 1923, Petrograd theorist Yuri Krutkov was going to Germany on a scientific visit. Friedmann discussed the essence of his arguments with Krutkov in the letter to Einstein and requested Krutkov to meet with Einstein and convince him that Friedmann was right. Krutkov met Einstein at Paul Ehrenfest house (both Krutkov and Einstein were Ehrenfest’s good friends) and Krutkov convinced Einstein that Friedmann was right. After that Einstein wrote a short note where he mentioned that ”My criticism, as I saw from Friedmann’s letter, communicated to me by Mr. Krutkov, was based on an error in calculations. I consider Friedman’s results to be correct and shed new light…” [33] [^6] Russian translation of the Einstein note published in https://ufn.ru/ufn63/ufn63_7/Russian/r637h.pdf. In 1923 Friedmann published his book “World as space and time” (it was the first popular presentation of GR in Russia). The cover page of the book is presented in Fig. 1.
@@
In 1924 Friedmann published his second cosmological paper, where the negative curvature case for the cosmological solution was considered [34] (English traslation of the paper was published in General Relativity and Gravitation [35]). [^7] According to Fock’s memoirs, he suggested Friedmann to consider this case when Fock translated the first cosmological Friedmann’s paper into German. Friedmann and Frederiks planned to write voluminous book ”Foundations of relativity theory” for Russian readers but the authors wrote only the first chapter ”Tensor calculus” which was published in 1924 and the project was not finished due to Friedmann’s untimely death. Investigations of atmosphere turbulence were among main Friedmann’s interests and in July 1925 he and P. F. Fedoseenko [^8] Pavel Fedoseenko was the Osoaviakhim-1 captain and died in the balloon crash on January 30, 1934. made a record-breaking balloon ascent (flight at 7400 m) where the flyers did medical experiments and measurements of atmosphere properties at different heights. After the flight Fedoseenko and Friedmann wrote their reports on the flight. In July – August 1925 Alexander Friedmann and his wife (N. E. Malinina – Friedmann) relaxed at the Crimea cost. On August 17, 1925 Alexander Friedmann came back to Leningrad, while his wife went to another town where she was supposed to work at the geomagnetic station. Perhaps, in his last train trip Friedmann was infected by a typhus since he ate dirty fruits. According to Friedmann himself, he probably got infected by eating an unwashed pear bought at one of the railway stations on the way from Crimea to Leningrad. Suddenly (on September 2) he felt sick. On September 16, 1925 Friedmann died in hospital (in days of the Baldin Seminar scientific community the gravitational community remembered this remarkable man and his works on cosmology, in which it was firstly discovered that the universe can expand). He was buried at the Smolensk Orthodox Cemetery in Leningrad (his tomb is presented in Fig. 2). As we discuss below models of evolving Universe were practically banned in USSR in 1933 – 1963 since the birth of the Universe may be easily associated with the Genesis in Bible text (we will discuss the issue below in the Chapter on Lemaître studies). In connection with the 75th anniversary of the birth of A. Friedmann, the Soviet authorities lifted the tacit ban on mentioning Friedmann’s cosmological works, and in June 1963, the physical-mathematical division of the Academy of Sciences organized a special session dedicated to his memory. At his introductory speech P. L. Kapitsa said at the session in June 1963 [36]:”Friedmann’s name has so far been undeservedly forgotten. This is unfair and it needs to be fixed. We must perpetuate this name. After all, Friedmann is one of the pioneers of Soviet physics, a scientist who made a great contribution to domestic and world science.” Soon after the session in July 1963 Soviet Physics Uspekhi published a special issue with articles of outstanding scientists on subjects which were connected with Friedman’s scientific interests. In his short note about a genesis of GR studies in Russia V.A. Fock wrote [37]: ”Alexander Alexandrovich Friedmann and Vsevolod Konstantinovich Frederiks were professors of Petrograd University (now Leningrad University) and they were the first scientists who started to teach Russian physicists (working in Petrograd) GR…Main speakers were V.K. Frederiks and A.A. Friedmann. Styles of their talks were rather different. Frederiks clearly understood physical aspects of theory and disliked to show mathematical calculations, while Friedmann placed emphasis on mathematics but not on physics. He aimed at mathematical rigor and paid attention to complete and precise formulations of initial assumptions. Discussions between Frederiks and Friedmann were very interesting.” At the session in 1963 it was decided to publish selected works of Friedmann as an academic publication edition [38] (it was very convenient for Soviet researchers and students to read Russian translations of fundamental Friedmann’s papers instead of difficult-to-access original articles written in German). In June 1988 Soviet Academy of Sciences organized a big international Friedmann seminar in Leningrad to celebrate the 100th anniversary of his birth.

### v3

@@ Abstract
NICA is a tool to investigate the early Universe in laboratory. It would be reasonable to recall some essential issues in physical cosmology development. GR was created by A. Einstein in 1915. In 1917 Einstein proposed the first (static) cosmological model. Soon after the A. Eddington proved that the model is unstable therefore it can not be realizable in nature. In 1922 and 1924 A. A. Friedmann found non-stationary solutions for cosmological equations written in the framework of GR. In 1927 G. Lemaître obtained very similar results and, in addition, he derived the Hubble law (E. Hubble obtained this law from observations). Unfortunately, G. Lemaître published his paper in not very popular Belgium journal. In 1931 Lemaître proposed the first version of hot Universe model (he called it hypothesis of the primeval atom). In his book «The Primeval Atom – An Essay on Cosmogony» Lemaître predicted even a background radiation as a signature of his model. At the end of 1940s G. Gamow and his students proposed his hot universe model where he explained primordial nucleosythesis of helium at the initial stage of the Universe evolution. One of the properties of Lemaître – Gamow model was an existence of CMB radiation with a temperature around a few K. It was recalled that the discovery of CMB radiation was done by T. Shmaonov in 1956 and his paper was published in 1957 (several years before the discovery done by Penzias and Wilson). In 1965, 1970 E. B. Gliner proposed vacuum like equation of matter which could correspond to exponential explosion of the Universe which was later called inflation. For decades in USSR, A. A. Friedmann’s cosmologocal non-stationary models were treated as purely mathematical results without cosmological and astronomical applications. On September 16, 1925 Friedmann passed away untimely and it would be reasonable to remind today his great contribution in physical cosmology since the authors book on Friedmann wrote “Similarly to Copernicus who forced the Earth to move, Friedmann forced the Universe to expand”.
[…]
@@ 1 Introduction
General Relativity (GR) was created after intensive conversations between A. Einstein and D. Hilbert in November 1915 [1, 2, 3]. Using his new theory of gravity in November 1915 A. Einstein explained the problem of the anomaly of Mercury’s motion, discovered in the middle of the 19th century by Le Verrier and remained unexplained for several decades. Einstein also considered the problem of deflecting a ray of light in a gravitational field, and it turned out that the angle of deviation in general relativity exceeds the angle of deviation in the Newtonian theory of gravity by two times [4] (see, English translation in [5]).
In 1917 A. Einstein found a static cosmological solution assuming that the Universe is homogeneous and isotropic [6] (see English translation in [7]). However, the assumptions adopted in the first cosmological paper were very useful and despite the fact that assumptions about the homogeinity and isotropy of the universe do not look very plausible, such assumptions turn out to be applicable to the theoretical description of astronomical observations. The assumptions are still used in a large number of studies on cosmology and are often called the cosmological principle. To create a static cosmological solution and to compensate gravitational attraction A. Einstein introduced so-called $\Lambda$- term which corresponds to repulsion. In many textbooks and articles the authors repeated after G. Gamow that ’’he [A. Einstein] remarked that the introduction of the cosmological term was the biggest blunder he ever made in his life’’ [8] (see also a similar sentence in a popular article [9])). However, taking into account Gamow’s personality, in particular, his predilection for jokes and practical jokes, and the different social status of Gamow and Einstein, it seems very strange that Einstein chose Gamow as an interlocutor to whom Einstein would admit to the biggest blunder of his entire life (see also discussions in [10, 11]). Moreover, in 1998 it was discovered an accelerated expansion of the Universe and in order to explain this phenomenon, it is necessary to introduce a non-zero value of the lambda term into the Einstein equations. Otherwise, it is necessary to use a some generalization of the lambda term, which has been called dark energy since 1999 [12, 13].
When discussing Einstein’s static cosmological model, it should be noted that it is hardly applicable to describe the behavior of the universe because as a famous Russian mathematician V. I. Arnold repeated only stable mathematical solutions exist in physical world. Really, analyzing properties of the static cosmological model in 1930 A. Eddington ’’working in conjunction with Mr. G. C. McVittie’’ proved that the Einstein’s static cosmological solution is unstable [14].
@@ 2 Start of GR studies in Russia
In the early 20th century, despite significant achievements in some areas of physical science development in Russia, world leaders in physics worked in Europe. Young Russians who wanted to acquire scientific skills at the cutting edge came to Europe for training and internship. One of these young people interested in science was a representative of the elite of the Russian empire (his father and grandfather were governors in Russian empire), Vsevolod Konstantinovich Frederiks (13.04.1885 (Warsaw) –- 06.01.1944 (Gorky)). He played a key role in initiating research in Russia in the field of GR [15, 16]. In 1903 Vsevolod Frederiks finished gymnasium in Nizhnij Novgorod and decided to go to Switzerland due to its convenient location to visit other European countries. In 1907 he finished Geneva University with specialisation in physics and in 1909 he got PhD degree in physics under supervision of professor G. E. Guye (Geneve). Since 1911 Frederiks was an assistant under professor Woldemar Voigt [^1] In 1887 W. Voigt found transformations which were similar to Lorentz ones [17]. In addition, Voigt described group properties of his transformations as it was noted by H. Minkowski in his famous lecture delivered in 1908 [18] (see also discussions about a relation between Voigt and Lorentz transformations in [19, 20, 21, 22]). at the Theoretical division of Physics Institute at Göttingen. In 1914 the WWI began and Frederiks started to be a civil prisoner since he was a citizen of an enemy country. In 1910s famous mathematician David Hilbert was very interested in physics he was searching for an educated physicist to discuss actual physical issues. Frederiks began working as a Hilbert’s personal assistant, and Hilbert paid Frederiks a monetary reward from his personal funds. The management of the University of Göttingen did not approve this practice to contact with representative of an enemy country in University and refused to allow Fredericks to enter the University in response to Hilbert’s appeal. In 1918 Frederiks came back in Russia and for a short period he worked in Institute of Physics and Biophysics in Moscow. In 1919 he moved to Petrograd and was a senior physicist at State Optical Institute, a member of Atomic commission, an associate professor in Petrograd State University, a professor in Pedagogical Institute, in 1920 he was a lecturer at Polytechnic Institute. In 1921 Frederiks published the first Russian review on general relativity (GR) in Soviet Physics Uspekhi which was a leading physical journal in Russia [23]. In 1923 Frederiks was a senior physicist in Institute of Physics and Technology, the Institute was founded by A. F. Ioffe in 1918 and in 1924 it was named the Leningrad Institute of Physics and Technology (LIPT) or shortly the Ioffe Institute. V. K. Frederiks and A. A. Friedmann [^2] British expert in astronomy and history of science S. Mitton noted [24]: ”There are three ways to spell his surname: Fridman, Friedmann, and Friedman… Friedmann is the most common version in use today, largely in connection with cosmology. I [S. Mitton] use Friedman when referring to his publications because that is the version he used as author. A Bolshevik Government Decree issued 1917 December 23 eliminated double letters in Russian orthography. Typographers were required to enforce the reforms. The contemporary Russian literature invariably uses Friedman.” I [AZ] basically use Friedmann in the paper since this spelling was used in a majority of articles and books that I read before writing the paper including English translation of the remarkable book [25] written by the best experts in history of Russian cosmology. decided to wrote a joint book on foundations of GR but in 1924 it was published only the first chapter of the book. In the chapter only a tensor calculus was discussed (this project to write a joint book was not realized due to the untimely death of A. A. Friedmann). In Autumn 1927 Frederiks married Maria Dmitrievna Shostakovich (a sister of composer D. D. Shostakovich). In 1931 Frederiks was appointed as the Head of the Crystallization Laboratory in LPTI and in 1933 started to be the Head of anisotropic liquid Laboratory in LPTI. Frederiks studies in liquid crystals were rather successful, in particular in 1927 V. Frederiks and A. Repiewa discovered the phase transition in liquid crystals under action of magnetic field. Now this property of liquid crystals is used in operation LC screens in different electronic devices and it is called now the Frederiks transition [26]. In 1934 Frederiks got DSc degree without a formal defense, he was nominated by the LPTI Scientific Council as a candidate for corresponding member of Soviet Academy of Science but he was not elected. Due to very rapid development of contemporary physics in 1920s and 1930s it was necessary to write new textbooks and V. K. Frederiks was the co-editor with A. Р. Afanasiev of Course on General Physics (I. K. Kikoin, Yu. B. Khariton were among authors of chapters in the book). Later, the textbook was republished where only A. Р. Afanasiev was indicated since Frederiks was in detention at this time. On October 20, 1936 V. K. Frederiks (among many other Leningrad scientists) was arrested as a defendant in the Pulkovo case. On May 23, 1937 Frederiks sentenced to 10 years in prison camps (Taishetlag). On January 6, 1944 Frederiks died. Only in 1957 his relatives got an official document on his death, where is the dash was written in the place of death. Lydia Chukovskaya (whose husband Matvey Bronstein [^3] M. Bronstein was one of the leaders of Soviet theoretical physics [27] and he was among the pioneers in quantum gravity research [28]. In the quoted paper C. Rovelli also noted that an introduction of the graviton concept proposed by D. I. Blokhintsev and F. M. Galperin in 1934 was an important initial step in development of quantum gravity. was an another Pulkovo case defendant) received a similar document concerning a time and a place of the Bronstein death [29].
@@
There’s no doubt that Friedmann’s cosmological work laid the foundations of physical cosmology and is a source of pride for Russian and international science [30, 31]. For Friedmann’s centenary, which was celebrated in June 1988, a scientific biography was written [32] (this remarkable book was translated into English in 1993). In the book the authors wrote, ’’He [Friedmann] discards a centuries-old tradition, which a priori, prior to all experience, considered the Universe eternal and forever motionless. He is making a real scientific revolution. Just as Copernicus forced the Earth to revolve around the Sun, so Friedmann forced the Universe to expand’’. In the days of the Baldin seminar we celebrated a memory of this remarkable scientist. Alexander Alexandrovich Friedmann (Friedman) was born on 4(16).06.1888 in Sankt-Peterburg (SPb) and died on 16.09.1925 in Leningrad (Sankt-Peterburg (1703 – 1914), Petrograd (1914 – 1924), Leningrad (1924 – 1991), Sankt-Peterburg (1991 – till now)). In 1897 - 1906 Alexander Friedmann was a student at the Second SPb gymnasium, in 1905 he wrote the first mathematical paper (published in 1907). In 1906 - 1910 he was student at mathematical division of the faculty of physics and mathematics. In 1910 - 1913 he left in the SPb University for a preparation for a professor position. In 1914 – 1916 Friedmann joined the Russian army as a volunteer, he served in aviation units. In 1918 – 1920 he was a professor of mechanics department of the Perm University. In 1920 – 1924 he was a researcher at the Atomic Commission in the State Optical Institute in Petrograd. In 1920 – 1924 he was a professor at the Faculty of Physics and Mechanics of the Petrograd Polytechnic Institute. In 1920 – 1925 Friedmann was a senior physicist, head of mathematical bureau, scientific secretary and since February 1925 director of the Main Geophysical Observatory. In 1922 Friedmann published his first cosmological paper [33] (English translation of the paper was published in [34]). In this paper, Friedman showed a solution describing an expanding universe that satisfies Einstein’s equations. Einstein read the Friedmann’s paper and he published a short comment [35] (consisting of a few sentences and one expression) where he claimed:’’The results of the work on the non-stationary world contained in the mentioned work seem suspicious to me. In fact, it turns out that the solution indicated in it does not satisfy the field equations.’’ [^4] A Russian translation of the comment published in the special issue of Soviet Physics – Uspekhi dedicated to the 75th anniversary of Friedmann https://ufn.ru/ufn63/ufn63_7/Russian/r637g.pdf. Friedmann was extremely disappointed by this opinion of the GR creator. Friedmann consulted with Petrograd mathematicians and sent a lengthy letter to Einstein where Friedmann presented formulas confirming his conclusions. At the end of his letter Friedmann requested Einstein [25] ’’…Should you find the calculations presented in my letter correct, please be so kind as to inform the editors of the Zeitschrift fur Physik about it; perhaps in this case you will publish a correction to your statement or provide an opportunity for a portion of this letter to be printed.’’ Friedman did not receive Einstein’s reply to this letter, possibly because Einstein was on a long overseas voyage at the time. In 1923, Petrograd theorist Yuri Krutkov was going to Germany on a scientific visit. Friedmann discussed the essence of his arguments with Krutkov in the letter to Einstein and requested Krutkov to meet with Einstein and convince him that Friedmann was right. In Leiden Krutkov met Einstein at the Paul Ehrenfest house (both Krutkov and Einstein were Ehrenfest’s good friends) and Krutkov convinced Einstein that Friedmann was right. After that Einstein wrote a short note where he mentioned that ’’My criticism, as I saw from Friedmann’s letter, communicated to me by Mr. Krutkov, was based on an error in calculations. I consider Friedman’s results to be correct and shed new light…’’ [36] [^5] Russian translation of the Einstein note published in https://ufn.ru/ufn63/ufn63_7/Russian/r637h.pdf. In 1923 Friedmann published his book “World as space and time” (it was the first popular presentation of GR in Russia). The cover page of the book is presented in Fig. 1.
@@
In 1924 Friedmann published his second cosmological paper, where the negative curvature case for the cosmological solution was considered [37]. English traslation of the paper was published in General Relativity and Gravitation [38]. [^6] According to Fock’s memoirs, he suggested Friedmann to consider this case when Fock translated the first cosmological Friedmann’s paper into German. Friedmann and Frederiks planned to write voluminous book ’’Foundations of relativity theory’’ for Russian readers but the authors wrote only the first chapter ’’Tensor calculus’’ which was published in 1924. The project was not finished due to Friedmann’s untimely death. Investigations of atmosphere turbulence were among main Friedmann’s interests and in July 1925 he and P. F. Fedoseenko [^7] Pavel Fedoseenko was the Osoaviakhim-1 captain and died in the balloon crash on January 30, 1934. made a record-breaking balloon ascent (flight at 7400 m) where the flyers did medical experiments and measurements of atmosphere properties at different heights. After the flight Fedoseenko and Friedmann wrote their reports on their measurements. In July – August 1925 Alexander Friedmann and his wife (N. E. Malinina – Friedmann) relaxed at the Crimea cost. On August 17, 1925 Alexander Friedmann came back to Leningrad, while his wife went to another town where she was supposed to work at the geomagnetic station. Perhaps, in his last train trip Friedmann was infected by a typhus since he ate dirty fruits. According to Friedmann himself, he probably got infected by eating an unwashed pear bought at one of the railway stations on the way from Crimea to Leningrad. Suddenly (on September 2) he felt sick. On September 16, 1925 Friedmann died in hospital (in days of the Baldin Seminar scientific community the gravitational community remembered this remarkable man and his works on cosmology, in which it was firstly discovered that the universe can expand). He was buried at the Smolensk Orthodox Cemetery in Leningrad (his tomb is presented in Fig. 2). As we discuss below models of evolving Universe were practically banned in USSR in 1933 – 1963 since the birth of the Universe may be easily associated with the Genesis in Bible text (we will discuss the issue below in the Chapter on Lemaître studies). In connection with the 75th anniversary of the birth of A. Friedmann, the Soviet authorities lifted the tacit ban on mentioning Friedmann’s cosmological works, and in June 1963, the physical-mathematical division of the Academy of Sciences organized a special session dedicated to his memory. At his introductory speech P. L. Kapitsa said at the session in June 1963 [39]:’’Friedmann’s name has so far been undeservedly forgotten. This is unfair and it needs to be fixed. We must perpetuate this name. After all, Friedmann is one of the pioneers of Soviet physics, a scientist who made a great contribution to domestic and world science.’’ Soon after the session in July 1963 Soviet Physics Uspekhi published a special issue with articles of outstanding scientists on subjects which were connected with Friedman’s scientific interests. In his short note about a genesis of GR studies in Russia V.A. Fock wrote [40]: "Alexander Alexandrovich Friedmann and Vsevolod Konstantinovich Frederiks were professors of Petrograd University (now Leningrad University) and they were the first scientists who started to teach Russian physicists (working in Petrograd) GR…Main speakers were V.K. Frederiks and A.A. Friedmann. Styles of their talks were rather different. Frederiks clearly understood physical aspects of theory and disliked to show mathematical calculations, while Friedmann placed emphasis on mathematics but not on physics. He aimed at mathematical rigor and paid attention to complete and precise formulations of initial assumptions. Discussions between Frederiks and Friedmann were very interesting." At the session in 1963 it was decided to publish selected works of Friedmann as an academic publication edition [41] (it was very convenient for Soviet researchers and students to read Russian translations of fundamental Friedmann’s papers instead of difficult-to-access original articles written in German). In June 1988 Soviet Academy of Sciences organized a big international Friedmann seminar in Leningrad to celebrate the 100th anniversary of his birth.